OpenStudy (anonymous):

Find the zeros x^4 -5x^2 +4

OpenStudy (triciaal):

let u = x^2 u^2 - 5 u + 4 = 0 can you factor this?

OpenStudy (triciaal):

remember u = x^2 x = +1, -1, +2, -2